Find the slope and the y-intercept of each line whose equation is given.
y=-9/10x-5



Write a point-slope equation for the line with the given slope that contains the given point.
m= 7/2;(-3,4)


Answer by Maths68(1474)

Find the slope and the y-intercept of each line whose equation is given.
y=-9/10x-5
Equation of the line
y=mx+b
Where m is slope of the of line b is y-intercept
Given line equation
y=-9/10x-5
arrange above equation
y=(-9/10)x+(-5)
Compare above equation with Equation of the line
we have
Slope = m = -9/10 and Y-intercept = b=-5


Write a point-slope equation for the line with the given slope that contains the given point.
m= 7/2;(-3,4)
Use point-slope form of Equation of the line which is
m=(y2-y1) /(x2-x1)
Put the given value
7/2=(y-4)/(x-(-3))
7/2=(y-4)/(x+3)
7(x+3)=2(y-4)
7x+21=2y-8
7x=2y-8-21
7x=2y-29
-2y=-7x-29
Divide by -2 both sides
-2y/-2=(-7x-29)/-2
y=-7x/-2 -29/-2
y=(7/2)x + 29/2
Above equation is required equation of the line